# Online Iterative Reinforcement Learning from Human Feedback with General Preference Model

### v2

@@ 4.1 Setup
In the offline setting, our goal is to learn a good policy from a pre-collected dataset $\mathcal{D}_{\mathrm{off}}=\{(x_{i},a_{i}^{1},a_{i}^{2},y_{i})\}_{i=1}^{n}$ without further query with the oracle $\mathbb{P}$, where comparison sample is assumed to be independently collected as in Equation (1). We measure the suboptimality of the learned policy $\hat{\pi}^{1}$ by the gap between the Nash value and the best response value:
@@
where the KL-regularized function $J$ is defined in Equation (3). Similar to the reward-based framework (Ouyang et al.,, 2022), one natural approach is a two-staged method:
  • Construct an empirical preference model (reward model in the literature) by maximizing the following log-likelihood function:
@@
  • Solve the policy by plugging the learned preference model $\hat{P}$ into the Nash Equilibrium Oracle 2.
However, it is known that this framework typically leads to severe reward over-optimization issue (Gao et al.,, 2023), meaning that while the model is preferred by the learned $\hat{P}$, it may not achieve good performance under the evaluation of $P^{*}$. This is because, with finite $\mathcal{D}_{\mathrm{off}}$ drawn from some behavior policy, it is unlikely to provide an accurate estimation for all the prompt-response pairs. Therefore, imposing heavy optimization pressure toward $\hat{P}$ will push the model to exploit these unreliable estimations to chase for a high proxy metric, which typically leads to a worse performance under the ground truth $P^{*}$. See Figure 1 for an illustration of this phenomenon in terms of reward learning.
Figure 1: An illustration of the reward over-optimization issue adapted from Gao et al., (2023). Here the proxy reward is trained from the responses of $\pi_{0}$. Therefore, in the early stage, the proxy reward aligns well with the gold reward in terms of the in-distribution responses. As the reward gets higher, the distribution shift becomes larger, and since the training set is lacking in the coverage over these out-of-distribution responses, the proxy reward does not align with the gold reward in this stage.
4.2 Learning with Pessimism
The recent advances in the offline RL theory have demonstrated that the principle of pessimism with a conservative estimation is statistically efficient for offline learning across a diverse set of scenarios (Jin et al.,, 2021; Rashidinejad et al.,, 2021; Xie et al.,, 2021; Zanette et al.,, 2021; Zhong et al., 2022a, ; Cui and Du,, 2022; Xiong et al.,, 2022; Zhang et al.,, 2023). In this section, we connect the KL-reversed minimax game in Equation (3) with offline RL by pessimism via version space [^2] We also introduce another algorithm achieving pessimism via uncertainty bonus construction, see Appendix B.2..
@@ 4.2 Learning with Pessimism
We introduce our algorithm, Pessimistic Equilibrium Learning from Human Feedback (PELHF) in ALgorithm 1. Given an offline dataset $\mathcal{D}_{\mathrm{off}}$, we first obtain the maximum likelihood estimation (MLE) $\hat{P}$ by maximizing Equation (6). Rather than directly planning with this empirical $\hat{P}$, we form a version space $\widehat{\mathcal{P}}$ that contains $P^{*}\in\widehat{\mathcal{P}}$ with a high probability under a suitable choice of $\beta$, as we show in the following lemma.
Lemma 1.
[Proof] Under Assumption 1, with probability at least $1-\delta$, we have
For each policy $\pi^{1}$, we take the minimum preference function over $\widehat{\mathcal{P}}$ and the best responded $\pi^{2}$ as its conservative value estimation:
@@
$J(\pi_{1}^{*},\pi_{2}^{*})-J(\hat{\pi}^{1},\dagger)\leq 4\beta\sqrt{\frac{\mathcal{C}(\pi_{*}^{1},\pi_{D},\mathcal{P})}{n}}.$
@@
This theorem shows that the suboptimality gap depends on how the target $(\pi^{1}_{*},\pi^{2})$ is covered by the offline dataset, where $\pi^{2}$ is maximized over the policy set $\Pi$. This coverage coefficient resembles the unilateral coverage [^3] In Appendix B.3, we show that with an improved analysis, Algorithm 1 enjoys a refined coverage condition, similar to the coverage notion in Zhang et al., (2023). for Markov games (Cui and Du,, 2022; Zhong et al., 2022a, ). Then, a natural question is whether a good coverage condition ($\mathcal{C}(\pi_{*}^{1},\pi_{D},\mathcal{P})$ is small) is practical in the context of LLMs. Unfortunately, since the response is usually long in practice, the distribution shift between policies is also very large. We summarize some statistics here:
  • Along the way of the RLHF training, the average density ratio $\frac{\pi(a|x)}{\pi_{0}(a|x)}>\exp(25)$ as reported in Figure 13 of Bai et al., 2022a . See similar results of rejection sampling fine-tuning (Dong et al.,, 2023) and DPO (Rafailov et al.,, 2023).
  • For a case study, we use the Gemma-7B-it as the behavior policy to collect data for aligning Gemma-2B-it (Team et al.,, 2024) with 15k prompt from (Cui et al.,, 2023). Then, we calculate the average KL divergence between Gemma-7B-it and Gemma-2B-it as $456.4$.
Therefore, it is unlikely to expect that we can learn the optimal policy from a pre-collected dataset. This motivates us to consider the online setting, where we can further query the preference oracle during the training to continuously enrich the dataset thus enhancing our models.
@@ 5.1 Setup of Iterative RLHF
The major difference between the online and offline settings is that online algorithms can further query the preference oracle $P^{*}$ along the way of training. Since updating the LLMs is expensive, we consider the batch online setting for a sparse policy update. Specifically, for each batch $t\in[T]$,
  • we first update the policy pair $(\hat{\pi}_{t}^{1},\hat{\pi}_{t}^{2})$ based on the historical information collected so far;
  • we collect $m$ tuples: we sample a random prompt by $x_{t,i}\sim d_{0}$, collect two responses by $(a_{t,i}^{1},a_{t,i}^{2})\sim(\hat{\pi}_{t}^{1},\hat{\pi}_{t}^{2})$, and query the preference signal $y_{t,i}\sim\mathrm{Ber}(P^{*}(x_{t,i},a_{t,i}^{1},a_{t,i}^{2}))$.
Here the batch size $m$ is usually very large compared to the typically adopted mini-batch update. To distinguish this from the sequential online setting where we update policy after collecting a single preference pair, we refer to this learning paradigm as the iterative RLHF.
@@ 5.2 Learning with Exploration
Information Ratio and Eluder Coefficient. Distinct from the offline setting where we assume the coverage condition of a pre-collected dataset $\mathcal{D}_{\mathrm{off}}$, online exploration makes it possible to upper bound the suboptimality by the complexity of the function space. We leverage the notion of the eluder coefficient, which explicitly limits the generalization from the visited state-action distributions to the unseen part.
@@
For any two policy $\pi^{1},\pi^{2}$, we define the information ratio as
@@
The information ratio and eluder coefficient considered here have also been adopted in the literature (e.g., Wang et al.,, 2020; Gentile et al.,, 2022; Xie et al.,, 2022; Ye et al., 2023a, ; Agarwal et al.,, 2023). Essentially, the information ratio compares the out-of-sample error on the unseen data with the in-sample error measured on the historical data, and can be interpreted as the worst-case ratio between them (as we take sup over all possible $P\in\mathcal{P}$). Meanwhile, the eluder coefficient limits the extent to which we can be “surprised” by the new out-of-sample distributions, given the historical data collected so far. The uncertainty for the preference model aligns with the uncertainty for the BT model under boundedness conditions, which is illustrated in the following example. We defer the details to Appendix C.1.
@@
Suppose the reward function can be embedded into a $d$-dimensional vector space $\{r(x,a)=\langle\theta,\phi(x,a)\rangle:\theta\in\mathbb{R}^{d},\|\theta\|\leq B,\|\phi(x,a)\|\leq 1\}$. Then, if we define the covariance matrix as $\Sigma_{t}=\sum_{s=1}^{t-1}\mathbb{E}_{x\sim d_{0},a^{1}\sim\hat{\pi}_{s}^{1},a^{2}\sim\hat{\pi}_{s}^{2}}(\phi(x,a^{1})-\phi(x,a^{2}))^{\top}(\phi(x,a^{1})-\phi(x,a^{2}))+\lambda(1+e^{B})^{2}I.$
@@
We refer interested readers to Du et al., (2021); Zhong et al., 2022b ; Xie et al., (2022) for the extensive examples when $d(\mathcal{P},\lambda,T)$ can have a sub-linear dependency on $T$. We are now ready to present the algorithm for the online setting, as summarized in Algorithm 2. Specifically, for each iteration:
  • The main agent exploits the information contained in the data collected so far by computing the MLE $\hat{P}_{t}$ and solving the minimax game with respect to it to get $\hat{\pi}_{t}^{1}$;
  • The enhancer, however, aims to facilitate the main agent’s learning by maximizing the uncertainty relative to the $\hat{\pi}_{t}^{1}$;
  • Finally, we use the policy pair to collect $m$ preference pairs and query oracle $P^{*}$ to get the preference signals.
@@
3:     Exploitation with the main agent: compute the MLE $\hat{P}_{t}$ with $\ell_{\mathcal{D}_{1:t-1}}$ defined in Equation (6) and compute Nash equilibrium by calling the Nash equilibrium oracle 2:
@@
$\pi^{2}_{t}=\mathop{\mathrm{argmax}}_{\pi^{2}\in\Pi}\widetilde{\Gamma}^{m}_{t}(\lambda,\hat{\pi}_{t}^{1},\pi^{2}):=\mathop{\mathrm{argmax}}_{\pi^{2}\in\Pi}\sup_{P\in\mathcal{P}}\frac{|\mathbb{E}_{x\sim d_{0}}[P(x,\hat{\pi}_{t}^{1},\pi^{2})-\hat{P}_{t}(x,\hat{\pi}_{t}^{1},\pi^{2})]|}{\sqrt{\lambda+\frac{1}{m}\sum_{s=1}^{t-1}\sum_{j=1}^{m}(P(x_{s,j},a_{s,j}^{1},a_{s,j}^{2})-\hat{P}_{t}(x_{s,j},a_{s,j}^{1},a_{s,j}^{2}))^{2}}},$ (10)
@@
In particular, we choose to adopt an algorithmic structure that assigns the exploitation and exploration to the main agent and enhancer, separately. This choice turns out to be important when we move toward practical algorithms with reasonable approximations, as we detail in Section 6. We now present the main theoretical guarantee for Algorithm 2.
@@
[Proof] Under Assumption 1, for any $\epsilon>0$, if we set the total iterations as $T=\min\{n\in\mathbb{N}^{+}:n\geq 2d(\mathcal{P},\lambda,n)\}$, batch size as $m=18T\log(2T|\mathcal{P}|/\delta)/\epsilon^{2}$, $\beta=\sqrt{2T\log(2T|\mathcal{P}|/\delta)/m}$, and $\lambda=2T\log(2T|\mathcal{P}|/\delta)/m$ for Algorithm 2, then, with probability at least $1-\delta$, there exists a $t_{0}\in[T]$,
@@
The theorem states that with suitable hyper-parameter choices, after roughly $T$ iterations (up to log factors), we can find an $\epsilon$-approximate Nash policy $\hat{\pi}_{t_{0}}^{1}$ for the max-player. Here $T$ depends on the eluder coefficient that is intrinsic to the preference model and characterizes the complexity of the RLHF problem.
5.3 Key Ideas
In this subsection, we present a brief discussion of the key algorithmic ideas. In Algorithm 2, the main agent always exploits all the historical information and takes the best guess we can obtain so far by solving the minimax game with respect to the MLE $\hat{P}_{t}$. However, the success of this process necessitates that the online data collected at each iteration provides enough information. Intuitively, we should explore the direction where we are uncertain about so that we can gain more information. To this end, the enhancer aims to find the $\hat{\pi}_{t}^{2}$ to maximize the uncertainty relative to the $\hat{\pi}_{t}^{1}$. Then, according to the definition of the eluder coefficient, we know that
Since each term on the left-hand side is non-negative, there exists at least a $t_{0}\in[T]$ such that the value at $t_{0}$ is smaller or equal to the average value:
$\min\big{(}1,(\Gamma_{t}(\lambda,\hat{\pi}_{t}^{1},\hat{\pi}_{t}^{2}))^{2}\big{)}\leq d(\mathcal{P},\lambda,T)/T\leq 1/2,$
which further implies that $(\Gamma_{t}(\lambda,\hat{\pi}_{t}^{1},\hat{\pi}_{t}^{2}))^{2}\leq 1/2$. Recalling that the uncertainty bonus is essentially the worst-case ratio between the out-of-sample error (our learning target) and the in-sample error, we know that for $\hat{\pi}_{t_{0}}^{1}$, our target is close to the controllable in-sample error. Specifically, similar to the proof of Lemma 1, we can show that the in-sample error satisfies (with details in the Appendix C):
$\frac{1}{m}\sum_{s=1}^{t-1}\sum_{j=1}^{m}(P^{*}(x_{s,j},a_{s,j}^{1},a_{s,j}^{2})-\hat{P}_{t}(x_{s,j},a_{s,j}^{1},a_{s,j}^{2}))^{2}\lesssim\frac{\log(2T|\mathcal{P}|/\delta)}{m}.$
@@ 5.3 Key Ideas
where $\beta$ is the parameter defined in Theorem 2. This set is never empty because we can prove that both $\hat{\pi}_{t}^{1}$ and $\mathop{\mathrm{argmin}}_{\pi^{\prime}}J(\hat{\pi}_{t}^{1},\pi^{\prime})$ belong to $\Pi_{t}$. Intuitively speaking, maintaining a small KL divergence against $\hat{\pi}_{t}^{1}$ corresponds to exploiting the historical information, and maximizes the uncertainty relative to $\hat{\pi}_{t}^{1}$ leading to more information gain. The choice of $\Pi_{t}$ represents a refined trade-off between these two different goals, which allows that $\hat{\pi}_{t}^{2}$ also converges to $\pi_{*}$. The details are deferred to Appendix C.2.
6 Practical Implementation of Iterative RLHF
In this section, we discuss how to implement the proposed theoretical algorithms.
6.1 Practical Algorithmic Design
Main agent approximates Nash equilibrium oracle via self-play IPO. The first step is to approximate the information-theoretical oracle 2. Due to the separation between exploitation and exploration, we do not introduce optimism for the main agent. In particular, the main agent aims to solve the minimax game with respect to a learned preference model that satisfies $\hat{P}_{t}(x,a^{1},a^{2})=1-\hat{P}_{t}(x,a^{2},a^{1})$. Optimizing a fixed and known preference model has been studied in Munos et al., (2023); Calandriello et al., (2024) and the proposed algorithms can serve as a reasonable approximation of the oracle. Specifically, we adopt the self-play IPO by optimizing the following loss function:

### v3

@@ 3.1 Setup
In the offline setting, our goal is to learn a good policy from a pre-collected dataset $\mathcal{D}_{\mathrm{off}}=\{(x_{i},a_{i}^{1},a_{i}^{2},y_{i})\}_{i=1}^{n}$ without further query with the oracle $\mathbb{P}$, where comparison sample is assumed to be independently collected as in (1). We measure the suboptimality of the learned policy $\hat{\pi}^{1}$ by the gap between the Nash value and the best response value:
@@
where the KL-regularized function $J$ is defined in (3). Similar to the reward-based framework [50], one natural approach is a two-staged method: (1) Construct an empirical preference model (reward model in the literature) by maximizing the log-likelihood function:
@@
(2) Solve the policy by plugging the learned preference model $\hat{P}$ into the Nash Equilibrium Oracle 2. However, this framework typically leads to severe reward over-optimization issue [26], meaning that while the model is preferred by the learned $\hat{P}$, it may not achieve good performance under the evaluation of $P^{*}$. This is because, with finite $\mathcal{D}_{\mathrm{off}}$ drawn from some behavior policy, it is unlikely to provide an accurate estimation for all the prompt-response pairs. Therefore, imposing heavy optimization pressure toward $\hat{P}$ will push the model to exploit these unreliable estimations to chase for a high proxy metric, thus leading to a worse performance under the ground truth $P^{*}$.
@@
3.2 Learning with Pessimism
The recent advances in the offline RL theory have demonstrated that the principle of pessimism with a conservative estimation is statistically efficient for offline learning across a diverse set of scenarios [35, 54, 69, 79, 86, 17, 71, 84]. In this section, we connect the KL-reversed minimax game in (3) with offline RL by pessimism via version space [^2] We introduce another algorithm achieving pessimism via uncertainty bonus construction, see Appendix C.2..
We introduce our algorithm, Pessimistic Equilibrium Learning from Human Feedback (PELHF) in Algorithm 1. Given an offline dataset $\mathcal{D}_{\mathrm{off}}$, we first obtain the maximum likelihood estimation (MLE) $\hat{P}$ by maximizing (7). Rather than directly planning with this empirical $\hat{P}$, we form a version space $\widehat{\mathcal{P}}$ that contains $P^{*}\in\widehat{\mathcal{P}}$ with a high probability under a suitable choice of $\beta$, as we show in Lemma 1. For each policy $\pi^{1}$, we take the minimum preference function over $\widehat{\mathcal{P}}$ and the best responded $\pi^{2}$ as its conservative value estimation:
@@ 3.2 Learning with Pessimism
$J(\pi_{1}^{*},\pi_{2}^{*})-J(\hat{\pi}^{1},\dagger)\leq 4\beta\sqrt{\mathcal{C}(\pi_{*}^{1},\pi_{D},\mathcal{P})/n}.$
@@
This theorem shows that the suboptimality gap depends on how the target $(\pi^{1}_{*},\pi^{2})$ is covered by the offline dataset, where $\pi^{2}$ is maximized over the policy set $\Pi$. This coverage coefficient resembles the unilateral coverage [^3] In Appendix C.3, we show that with an improved analysis, Algorithm 1 enjoys a refined coverage condition, similar to the coverage notion in [84]. for Markov games [17, 86]. Then, a natural question is whether a good coverage condition ($\mathcal{C}(\pi_{*}^{1},\pi_{D},\mathcal{P})$ is small) is practical in the context of LLMs. Unfortunately, since the response is usually long in practice, the distribution shift between policies is also very large. We summarize some observations here. First, along the way of the RLHF training, the average density ratio $\pi(a|x)/\pi_{0}(a|x)>\exp(25)$ as reported in Figure 13 of Bai et al. [4]. See similar results of rejection sampling fine-tuning [18] and DPO [53]. Second, for a case study, we use the Gemma-7B-it as the behavior policy to collect data for aligning Gemma-2B-it [59] with 15k prompt from [16]. Then, we calculate the average KL divergence between Gemma-7B-it and Gemma-2B-it as $456.4$. This evidence indicates that the coverage coefficient probably explodes in realistic scenarios. Therefore, it is unlikely to expect that we can learn the optimal policy from a pre-collected dataset. This motivates us to consider the online setting, where we can further query the preference oracle during the training to enrich the dataset thus enhancing our models continuously.
@@ 4.1 Setup of Iterative RLHF
The major difference between the online and offline settings is that online algorithms can further query the preference oracle $P^{*}$ along the way of training. Since updating the LLMs is expensive, we consider the batch online setting for a sparse policy update. Specifically, for each batch $t\in[T]$, we first update the policy pair $(\hat{\pi}_{t}^{1},\hat{\pi}_{t}^{2})$ based on the historical information collected so far. Then, we collect $m$ tuples: we sample a random prompt by $x_{t,i}\sim d_{0}$, collect two responses by $(a_{t,i}^{1},a_{t,i}^{2})\sim(\hat{\pi}_{t}^{1},\hat{\pi}_{t}^{2})$, and query the preference signal $y_{t,i}\sim\mathrm{Ber}(P^{*}(x_{t,i},a_{t,i}^{1},a_{t,i}^{2}))$. Here the batch size $m$ is usually very large compared to the typically adopted mini-batch update. To distinguish this from the sequential online setting where we update policy after collecting a single preference pair, we refer to this learning paradigm as the iterative RLHF.
@@ 4.2 Learning with Exploration
Information Ratio and Eluder Coefficient. Distinct from the offline setting where we assume the coverage condition of a pre-collected dataset $\mathcal{D}_{\mathrm{off}}$, online exploration makes it possible to upper bound the suboptimality by the complexity of the function space. We leverage the notion of the eluder coefficient, which limits the generalization from visited state-action distributions to unseen parts.
@@
At round $t$, given an estimation $\hat{P}\in\mathcal{P}$, we define the information ratio for any two policy $\pi^{1},\pi^{2}$ as
@@
The information ratio and eluder coefficient considered here have also been adopted in the literature [e.g., 64, 28, 70, 74, 1]. Essentially, the information ratio compares the out-of-sample error on the unseen data with the in-sample error measured on the historical data, and can be interpreted as the worst-case ratio between them (as we take supreme over all possible $P\in\mathcal{P}$). Meanwhile, the eluder coefficient limits the extent to which we can be “surprised” by the new out-of-sample distributions, given the historical data collected so far. The uncertainty for the preference model aligns with the uncertainty for the BT model under boundedness conditions, which is illustrated in the following example. We defer the details to Appendix D.1.
@@
Suppose the reward function can be embedded into a $d$-dimensional vector space $\{r(x,a)=\langle\theta,\phi(x,a)\rangle:\theta\in\mathbb{R}^{d},\|\theta\|\leq B,\|\phi(x,a)\|\leq 1\}$. Then, if we define the covariance matrix as $\Sigma_{t}=\sum_{s=1}^{t-1}\mathbb{E}_{x\sim d_{0},a^{1}\sim\hat{\pi}_{s}^{1},a^{2}\sim\hat{\pi}_{s}^{2}}(\phi(x,a^{1})-\phi(x,a^{2}))^{\top}(\phi(x,a^{1})-\phi(x,a^{2}))+\lambda(1+e^{B})^{2}I,$ we have
@@
3:     Exploitation with the main agent: compute the MLE $\hat{P}_{t}$ with $\ell_{\mathcal{D}_{1:t-1}}$ defined in (7) and compute Nash equilibrium by calling the Nash equilibrium oracle 2:
@@
$\pi^{2}_{t}=\mathop{\mathrm{argmax}}_{\pi^{2}\in\Pi}\widetilde{\Gamma}^{m}_{t}(\lambda,\hat{\pi}_{t}^{1},\pi^{2}):=\sup_{P\in\mathcal{P}}\frac{|\mathbb{E}_{x\sim d_{0}}[P(x,\hat{\pi}_{t}^{1},\pi^{2})-\hat{P}_{t}(x,\hat{\pi}_{t}^{1},\pi^{2})]|}{\sqrt{\lambda+\frac{1}{m}\sum_{s=1}^{t-1}\sum_{j=1}^{m}(P(x_{s,j},a_{s,j}^{1},a_{s,j}^{2})-\hat{P}_{t}(x_{s,j},a_{s,j}^{1},a_{s,j}^{2}))^{2}}},$ (11)
@@
We refer interested readers to Du et al. [20], Zhong et al. [87], Xie et al. [70] for the extensive examples when $d(\mathcal{P},\lambda,T)$ can have a sub-linear dependency on $T$. We are now ready to present the algorithm for the online setting, as summarized in Algorithm 2. Specifically, for each iteration, the main agent exploits the information contained in the data collected so far by computing the MLE $\hat{P}_{t}$ and solving the minimax game with respect to it to get $\hat{\pi}_{t}^{1}$. The enhancer, however, aims to facilitate the main agent’s learning by maximizing the uncertainty relative to the $\hat{\pi}_{t}^{1}$. Finally, we use the policy pair to collect $m$ preference pairs and query oracle $P^{*}$ to get the preference signals. Notably, to facilitate the computation for the main agent, instead of adding optimism to the value function, we impose the exploration role on the enhancer. This choice turns out to be important when we move toward practical algorithms with reasonable approximations, as we detail in Section 5. We now present the main theoretical guarantee for Algorithm 2.
@@
[Proof] Under Assumption 1, for any $\epsilon>0$, if we set the total iterations $T=\min\{n\in\mathbb{N}^{+}:n\geq 2d(\mathcal{P},\lambda,n)\}$, batch size $m=18T\log(2T|\mathcal{P}|/\delta)/\epsilon^{2}$, $\beta=\sqrt{2T\log(2T|\mathcal{P}|/\delta)/m}$, and $\lambda=2T\log(2T|\mathcal{P}|/\delta)/m$ for Algorithm 2, then, with probability at least $1-\delta$, there exists a $t_{0}\in[T]$,
@@
The theorem states that with suitable hyper-parameter choices, after $T$ iterations (up to log factors), we can find an $\epsilon$-approximate Nash policy $\hat{\pi}_{t_{0}}^{1}$ for the max-player. Here $T$ depends on the eluder coefficient that is intrinsic to the preference model and characterizes the complexity of the problem.
Key Ideas. We present a brief discussion of the key analysis ideas. Similar to Lemma 1, the MLE $\hat{P}$ ensures a controllable in-sample error (with details in the Appendix D). Recalling that the uncertainty bonus is essentially the worst-case ratio between the out-of-sample error (our learning target) and the in-sample error, to finally bound the out-of-sample error, we need to explore each direction where we are uncertain about so that the average uncertainty bonus is sufficiently small. Since the main agent is greedy (takes the best guess we can obtain so far), the enhancer plays the exploration role by maximizing the uncertainty relative to the $\hat{\pi}_{t}^{1}$. Then, since the eluder dimension is finite: $\sum_{t=1}^{T}\min\big{(}1,(\Gamma_{t}(\lambda,\hat{\pi}_{t}^{1},\hat{\pi}_{t}^{2}))^{2}\big{)}\leq d(\mathcal{P},\lambda,T)$, there exists at least a $t_{0}\in[T]$ such that the value at $t_{0}$ is smaller or equal to the average value:
$\min\big{(}1,(\Gamma_{t}(\lambda,\hat{\pi}_{t}^{1},\hat{\pi}_{t}^{2}))^{2}\big{)}\leq d(\mathcal{P},\lambda,T)/T\leq 1/2.$
Hence, with a proper $m$, we can obtain the result of Theorem 2.
@@
where $\beta$ is the parameter defined in Theorem 2. This set is never empty because we can prove that both $\hat{\pi}_{t}^{1}$ and $\mathop{\mathrm{argmin}}_{\pi^{\prime}}J(\hat{\pi}_{t}^{1},\pi^{\prime})$ belong to $\Pi_{t}$. Intuitively, maintaining a small KL divergence against $\hat{\pi}_{t}^{1}$ corresponds to exploiting the historical information, and maximizing the uncertainty relative to $\hat{\pi}_{t}^{1}$ leads to more information gain. The choice of $\Pi_{t}$ represents a refined trade-off between these two different goals, thus making $\hat{\pi}_{t}^{2}$ also converge to $\pi_{*}$. The details are deferred to Appendix D.2.
5 Practical Implementation of Preference Model and Iterative RLHF
In this section, we discuss how to implement the theoretical Algorithm 2 for the online setting.
Main agent approximates Nash equilibrium oracle via self-play IPO. Approximating the information-theoretical oracle 2 given a known preference model has been studied in Munos et al. [46], Calandriello et al. [11]. The proposed algorithm, self-play IPO, can serve as a reasonable approximation of the oracle by optimizing the following loss function: